OpenStudy (anonymous):

solve -3x + 1 + 5x = 7 +15 + 7x

OpenStudy (anonymous):

add like terms

OpenStudy (anonymous):

-5X=11 X=-11/5

OpenStudy (anonymous):

\[-3x+1+5x=7+15+7x\] \[2x+1=7x+22\] \[1-22=7x-2x\] \[-21=5x\] \[-\frac{21}{5}=x\]

OpenStudy (anonymous):

not sure where the -11 came from